# Mastering NIM and Impartial Games with Weak Neural Networks: An AlphaZero-inspired Multi-Frame Approach

### v2

@@ Abstract
We introduce a practical circuit-complexity model for fixed-precision neural networks and apply it to explain (and partially overcome) a persistent learnability barrier in impartial games, with NIM as the canonical example. Our model captures bounded-depth, polynomial-size, fixed-precision neural inference (including finite-window recurrent and attention-style models) and shows that these models can be simulated by $\mathrm{AC}^{0}$ circuits. This places them below $\mathrm{TC}^{0}$ and, in particular, below exact parity and exact majority capabilities.
On the negative side, we prove that single-frame AlphaZero-style agents with such $\mathrm{AC}^{0}$-constrained evaluation/policy networks cannot achieve strong mastery of NIM (optimal play from every reachable position), even when augmented with any polynomial-time bounded search. The core obstruction is the necessity of global parity (nim-sum) information.
On the positive side, we show that augmenting the state with a short window of recent history (two frames suffice for NIM) exposes locally computable transition information. In particular, nimber differences between consecutive positions are $\mathrm{AC}^{0}$-computable when the positions differ in a constant number of heaps. This supports a simple local restoration response rule: after an opponent move, one can restore the zero nim-sum invariant by matching the observed nimber difference, without ever recomputing the global nim-sum from scratch.
Empirically, we validate the restoration viewpoint in a controlled supervised setting: from teacher–random transitions in 20-heap, 4-bit NIM, the two-frame policy exhibits a two-stage learning curve (an intermediate plateau around 0.54 followed by near-perfect restoration accuracy), while the one-frame baseline stays near chance. When combined with bounded per-move MCTS, the learned 2F priors translate into a large head-to-head advantage (Section 6.2).
Finally, we address an important practical question: why $\mathrm{AC}^{0}$ is a good model for feasible learnability even though practitioners observe that neural networks can learn “majority-like” behaviour. The resolution is that both $\mathrm{AC}^{0}$ and realistic fixed-precision networks can learn and compute approximate majority very well, while the sharpness and reliability needed to leverage majority gates to compute parity (as in $\mathrm{TC}^{0}$ constructions) is not feasible under fixed precision and noisy training/inference.
@@ 1 Introduction
AlphaZero-style reinforcement learning has achieved striking successes across a range of perfect-information games [18, 17, 15, 16]. Motivated by Harvey Friedman’s proposal of NIM as an AI challenge [8], a series of empirical studies explored whether “standard” AlphaZero pipelines can master impartial games such as NIM [24, 22, 25]. A recurring phenomenon is that, despite the apparent simplicity of NIM, learning optimal play is blocked by global parity constraints (the nim-sum).
This paper develops a theory that both (i) explains why this barrier is robust under standard single-frame training and (ii) shows how a representation change (multi-frame history) enables optimal play without requiring more powerful neural computation.
@@ 1.1 Contributions
  • An $\mathrm{AC}^{0}$ model for fixed-precision neural inference. We formalize three “weak” neural models (feedforward threshold networks, finite-window recurrent networks, and finite-window attention-style models) with constant depth, polynomial size, and fixed precision weights/thresholds. We show that they can be simulated by $\mathrm{AC}^{0}$ circuits (Theorem 2.2), with a clean construction in Appendix B.
  • Single-frame impossibility for strong mastery. We prove that $\mathrm{AC}^{0}$-constrained single-frame policy/evaluation networks cannot strongly master NIM, and polynomial-time bounded search cannot overcome this (Section 4.2).
  • Multi-frame possibility via local delta and restoration. With two-frame (or bounded) history, we show that local nimber-difference computations are $\mathrm{AC}^{0}$-computable (Lemma 4.5) and that a player can restore the zero nim-sum invariant after an opponent move by matching the observed nimber difference (Proposition 4.6). This bypasses the need to compute the global nim-sum from scratch on every move.
  • Empirical validation via supervised restoration and search. We instantiate the restoration principle as a supervised learning problem using teacher–random transitions: the input is either the current position (1F) or the last two positions (2F), and the label is a (possibly non-unique) optimal restoration/winning move from the post-opponent position. In a 20-heap, 4-bit NIM setting with a dataset of $10^{6}$ examples, the one-frame model remains at a near-random baseline (about $8\%$ restoration accuracy), while the two-frame model exhibits a two-stage learning curve: it first rises to an intermediate plateau around $0.54$ and later breaks through to essentially perfect restoration accuracy. When paired with per-move MCTS (fixed simulation budget), the learned 2F priors yield a large head-to-head Elo advantage over 1F under the same search budget (Section 6.2).
  • Why $\mathrm{AC}^{0}$ is a good practical learnability model. We clarify the apparent tension that “majority feels feasible” in practice while parity remains infeasible: what is learnable/feasible is approximate majority, which is compatible with $\mathrm{AC}^{0}$, whereas using majority to compute parity requires a level of sharpness/precision that is not feasible under fixed precision and noise (Section 5).
@@ 1.2 Paper organisation
Section 2 introduces the fixed-precision neural models and the $\mathrm{AC}^{0}$ simulation theorem. Section 3 recalls impartial games and NIM. Section 4 presents the impossibility and possibility results. Section 5 discusses learnability and the $\mathrm{AC}^{0}$ vs $\mathrm{TC}^{0}$ modelling issue. Section 6 reports experiments, including a supervised restoration benchmark that sharply separates one-frame and two-frame policies (Section 6.2). Appendices provide formal definitions and full circuit constructions.
@@ 2 Theoretical framework and neural network models
2.1 Fixed-precision neural models and $\mathrm{AC}^{0}$ simulation
We consider neural inference models that are (i) bounded depth, (ii) polynomial size, and (iii) fixed precision: all weights and thresholds are drawn from a finite set of representable values that does not grow with the input length. Concretely, we fix constants:
  • a precision parameter (e.g. bit-width) captured by a fixed denominator $D\in\mathbb{N}$,
  • a magnitude bound $W\in\mathbb{N}$,
@@ 2.1 Fixed-precision neural models and $\mathrm{AC}^{0}$ simulation
$\mathcal{Q}_{W,D}\;=\;\Big\{\frac{k}{D}\ :\ k\in\{0,1,2,\dots,W\}\Big\},$
@@
Definition 2.1 (Neural networks with constant precision (informal)).
A fixed-precision $\mathsf{NN}$ is a constant-depth feedforward network of polynomially many threshold neurons, whose weights and thresholds lie in $\mathcal{Q}_{W,D}$. Analogously, fixed-precision $\mathsf{RNN}$ and $\mathsf{LTST}$ models have constant depth and polynomial size per time step and operate within a fixed time window $T$.
Why $\mathrm{AC}^{0}$ and not $\mathrm{TC}^{0}$? (Key modelling paragraph)
A common first reaction is to model constant-depth threshold networks by $\mathrm{TC}^{0}$, since $\mathrm{TC}^{0}$ includes majority/threshold gates. However, the definition of $\mathrm{TC}^{0}$ (as used in circuit complexity) implicitly allows thresholds and weights whose description length can scale with input length (typically $\Theta(\log n)$ bits, or more generally polylogarithmic), enabling exact majority and other sharp global counting predicates. In contrast, our focus is fixed precision neural inference: every real parameter is stored in a fixed word size (e.g. 32-bit), so the set of representable thresholds and weights is finite and does not scale with $n$. Under this constraint, the “knife-edge” sharpness required for exact majority on arbitrarily large inputs is not available; what is feasible in practice is learning approximate majority behaviour (high accuracy away from the threshold), which is already achievable within $\mathrm{AC}^{0}$. Crucially, the degree of sharpness needed to use majority gates as reliable building blocks for parity computations is not feasible under fixed precision and noise; this explains why parity remains unlearnable/unstable even though majority-like behaviour is learnable (expanded in Section 5).
@@ 2.2 Simulation by circuits
Theorem 2.2 (Simulation of fixed-precision neural models by $\mathrm{AC}^{0}$).
@@
    (Fixed precision) all weights and thresholds lie in $\mathcal{Q}_{W,D}$.
@@ 3 Impartial games and NIM
We briefly recall standard definitions; see [2, 3, 4, 5] for extensive background.
@@ 3.2 NIM
A NIM position consists of $h$ heaps with sizes $x_{1},\dots,x_{h}\in\mathbb{N}$. A legal move reduces exactly one heap: choose $i$ with $x_{i}>0$ and replace $x_{i}$ by some $x_{i}^{\prime}<x_{i}$.
@@ 4.1 Mastery notions
We distinguish between two goals:
Definition 4.1 (Strong vs. weak mastery).
An agent achieves strong mastery of NIM if it plays optimally from every reachable position (equivalently, implements a winning strategy whenever one exists). It achieves weak mastery relative to an initial position $p_{0}$ if it plays optimally on the game tree reachable from $p_{0}$.
@@ 4.2 Single-frame impossibility results
Theorem 4.2 (Single-frame strong mastery requires parity; impossible in $\mathrm{AC}^{0}$).
Fix any encoding of NIM positions into Boolean strings of length $n$ such that each heap size is represented in binary with $\Theta(\log m)$ bits (so XOR of heap sizes corresponds to a parity-type computation over $\Theta(n)$ bits). No family of fixed-precision $\mathsf{NN}/\mathsf{RNN}/\mathsf{LTST}$ models (Definition 2.2) using only a single-frame input can achieve strong mastery of NIM for unbounded input sizes.
Proof sketch.
By Theorem 2.2, any such model can be simulated by $\mathrm{AC}^{0}$. Strong mastery implies that from every position the agent can decide (implicitly, via action choice) whether the position is winning or losing, which in NIM is equivalent to whether the global XOR (nim-sum) is zero (Theorem 3.4). This yields an $\mathrm{AC}^{0}$ computation of a parity-type predicate on $\Theta(n)$ bits, contradicting $\mathrm{PARITY}\notin\mathrm{AC}^{0}$ [9, 11, 19, 13]. ∎
Theorem 4.3 (Polynomial-time bounded search does not fix single-frame $\mathrm{AC}^{0}$ limitations).
Consider an AlphaZero-style agent that, at each node, uses fixed-precision single-frame $\mathrm{AC}^{0}$-constrained policy/value networks and performs any polynomial-time bounded tree search (including MCTS variants) whose node expansions and back-ups are polynomially bounded in the input length. Then such an agent still cannot achieve strong mastery of NIM for unbounded input sizes.
Proof sketch.
Intuitively, MCTS acts as a breadth-increaser: it allocates its simulation budget according to a policy prior and/or a learned value signal. If these signals cannot reliably distinguish nim-sum parity, then (in the regime of large heaps) the search remains essentially blind to the winning manifold: it has no robust criterion for concentrating visits on the unique restoring reductions rather than on the exponentially many non-restoring moves.
If the evaluation/policy signals are $\mathrm{AC}^{0}$-bounded and cannot reliably encode nim-sum parity, then the search lacks a systematic advantage in prioritizing winning moves as the branching factor grows. In NIM with large heaps (binary encoding), the fraction of winning moves can become extremely small in worst cases, and without parity information the search cannot consistently prune or focus to guarantee optimality within polynomial effort. A full formalization can be phrased as a reduction: any such search procedure that succeeds would induce (via transcript-to-circuit transformation) a polynomial-size constant-depth circuit family deciding nim-sum parity, again contradicting $\mathrm{PARITY}\notin\mathrm{AC}^{0}$. ∎
4.2.1 Implications for weak mastery
The above impossibility is for strong mastery. Weak mastery can sometimes be achieved from special initial positions with strong symmetry. For example, positions consisting of paired equal heaps permit mirroring strategies (Appendix D).

### v3

@@ Abstract
[…]
Keywords: impartial games; NIM; AlphaZero; circuit complexity; quantised inference; fixed latency; planning; Monte Carlo tree search.
@@ 1 Introduction
Game play has historically served as the “Formula 1” of artificial intelligence: a high-speed, controlled proving ground where engineering limits are tested and new paradigms are born. From Deep Blue in chess to AlphaGo in Go, games have repeatedly driven advances in search, planning, and reinforcement learning [21, 23, 22, 20]. Negative results have been equally influential: Minsky and Papert’s impossibility result for single-layer perceptrons on parity/XOR redirected the field towards multilayer architectures [11].
History demonstrates that optimal algorithms are often reverse-engineered from such lower bounds: Nesterov’s accelerated method and worst-case optimal join algorithms both emerged from explicit complexity limits [15, 16]. We adopt this methodology by using circuit complexity not merely to predict failure, but to derive the minimal architectural priors required for scalable and robust reasoning. Our contribution is an explicit barrier $\to$ bypass methodology: each barrier we prove is paired with a targeted representational or hybrid step motivated by the barrier mechanism.
Today, despite the success of AlphaZero-style methods, such a barrier appears in impartial games such as NIM. Motivated by Harvey Friedman’s proposal of NIM as an AI challenge [4], several empirical studies asked whether standard AlphaZero pipelines can master this setting [31, 29, 32]. A recurring phenomenon is that, despite NIM’s simple rules, learning optimal play is blocked by a global parity invariant (the nim-sum).
This paper formalises that barrier using circuit complexity and derives the structural bypasses. The central observation is that while the state can be hard to process, the transition can be easy: augmenting with short history converts a global hardness obstacle into a local verification task. Figure 1 summarises this viewpoint for NIM.
Input $P_{t}$ (Global) Single-Head Network $\times$ Cannot compute
global parity
[…]
(a) Barrier: Single-Frame, Single-Head Input $(P_{t-1},P_{t})$ (Transition) Single-Head Network $\checkmark$ Computes Local
Difference $\Delta$
(b) Bypass 1: Multi-Frame Input Input $P_{t}$ (Global) Head $\pi_{0}$ Head $\pi_{k}$ $\vdots\checkmark$ Heads recover
$B$ parity bits
(c) Bypass 2: Multi-Head Output
[…]
@@ 1.1 Contributions
  • $\mathrm{AC}^{0}$ containment for fixed-scale quantised inference. We formalise three families of bounded-latency neural models (feedforward threshold networks, finite-window recurrent networks, and finite-window attention-style models) operating in a fixed-scale quantised regime (bounded range and fixed quantisation grid independent of input length), and show they are simulable by $\mathrm{AC}^{0}$ circuits (Theorem 2.4).
  • Single-frame impossibility for strong mastery. Under this model, no single-frame policy/value network can strongly master NIM in the worst case (Theorem 5.2).
  • Planning analysis and the single-policy bottleneck. We analyse MCTS-style rollouts as an interface between a weak evaluator and a long-horizon game. With $B$ distinct rollout heads one can recover all $B$ bits of the nim-sum by depth amplification (Proposition 5.9), whereas a single rollout head reveals at most one fixed linear functional of the nim-sum regardless of search budget (Proposition 5.10 and Theorem 5.14).
  • Two-frame bypass via local differences and restoration. With two consecutive frames, the agent can compute a local nimber difference $\Delta(P,P^{\prime})$ in $\mathrm{AC}^{0}$ (Lemma 5.4) and implement a restoration rule that maintains $\operatorname{nim}=0$ (Proposition 5.5), yielding a universal-verifier rollout policy (Proposition 5.13).
  • Empirical validation and learnability discussion. On 20-heap, 4-bit NIM with $10^{6}$ supervised restoration examples, a one-frame model stays near chance while a two-frame model learns restoration (Section 7.2). In a complementary single-frame experiment, multi-head FSM-controlled shootouts achieve perfect win/loss position classification while the single-head baseline remains bottlenecked by one-bit information (Section 7.3). We also discuss why $\mathrm{AC}^{0}$ is a useful abstraction for what is robustly feasible under fixed-scale quantised inference (Section 6.3).
1.2 Scope and limitations
Our theorems are worst-case representational statements about what can be computed under explicit architectural constraints. They are not training-complexity results, and they intentionally abstract away from several practical details.
  • Fixed-scale quantisation / fixed latency. The $\mathrm{AC}^{0}$ simulation (Theorem 2.4) assumes constant depth and a fixed-scale, bounded-range quantisation regime: weights/thresholds lie on a fixed grid whose range/resolution do not scale with the input length (cf. Remark 2.2). Architectures that rely on fan-in dependent re-normalisation, biases/thresholds growing with dimension, or depth growing with $n$ fall outside this abstraction.
  • Theoretical idealisation vs. practical normalisation layers. Our formal abstraction targets the logical reasoning layers that must compute parity-type invariants. The experimental networks also include LayerNorm and Softmax (implemented as Log-Softmax), which are statistical normalisation layers with global reductions and therefore exceed strict $\mathrm{AC}^{0}$ as literal circuit primitives. We treat these layers as calibration/normalisation operations, distinct from the logical reasoning layers, rather than as sources of parity-computing power. Empirically, this distinction is consistent with Section 7.2: the one-frame model still fails despite having the same normalisation layers.
  • Stylised planning interface. Our rollout results model planning as querying deterministic rollout outcomes under a specified policy head. This captures the information that a weak evaluator can expose through repeated simulation, but it does not model all details of practical MCTS (stochastic policies, exploration noise, learned value backup, etc.).
  • No guarantee of learnability. The multi-policy-head construction proves existence of a strong-mastery algorithm under multiple heads; we do not claim that self-play efficiently discovers the required structured heads. Our experiments therefore focus on a controlled supervised restoration signal that isolates the representational difference between one-frame and two-frame inputs.
@@ 1.3 Paper organisation
Section 2 introduces the fixed-scale quantised inference model and the $\mathrm{AC}^{0}$ simulation theorem. Section 3 situates our contribution in circuit complexity, modern sequence models, and reinforcement learning for combinatorial games. Section 4 recalls NIM and the nim-sum characterisation. Section 5 presents the main impossibility and bypass results, including the rollout analysis in Section 5.4. Section 6 discusses modelling choices and learnability considerations. Section 7 reports experiments. Appendices contain formal model definitions and proofs.
@@ 2 Theoretical framework and neural network models
2.1 Fixed-scale quantised inference and $\mathrm{AC}^{0}$ simulation
We consider neural inference models that are (i) bounded depth, (ii) polynomial size, and (iii) operate in a fixed-scale quantised regime: all weights and thresholds lie on a fixed quantisation grid of bounded dynamic range that does not change with the input length. This is a stylised abstraction of deployment-time low-bit inference (e.g. INT8/fixed-point) with fixed scaling factors, where one does not re-normalise weights by fan-in as the problem size grows.
Concretely, we fix constants:
  • a quantisation resolution parameter captured by a fixed denominator $D\in\mathbb{N}$ (grid step $1/D$),
  • a dynamic-range bound $W\in\mathbb{N}$ (maximum magnitude $W/D$),
@@ 2.1 Fixed-scale quantised inference and $\mathrm{AC}^{0}$ simulation
$\mathcal{Q}_{W,D}\;=\;\Big\{\frac{\ell}{D}\ :\ \ell\in\{0,1,2,\dots,W\}\Big\},$
@@
Definition 2.1 (Neural networks with fixed-scale quantised parameters (informal)).
A fixed-scale quantised $\mathsf{NN}$ is a constant-depth feedforward network of polynomially many threshold neurons, whose weights and thresholds lie in $\mathcal{Q}_{W,D}$. Analogously, fixed-scale quantised $\mathsf{RNN}$ and $\mathsf{LTST}$ models have constant depth and polynomial size per time step and operate within a fixed time window $T$.
Why fixed-scale quantisation suggests an $\mathrm{AC}^{0}$ abstraction.
The key structural property is fixed scale: the smallest nonzero weight is $1/D$ and the largest threshold is $W/D$, with $W,D$ independent of the input length. Consequently, whenever a threshold unit outputs $1$, there is a satisfying witness consisting of at most $W$ input literals whose contributions already cross the threshold. This yields a constant-width DNF expansion for each unit (Lemma B.1), and composing a constant number of layers yields an $\mathrm{AC}^{0}$ simulation (Theorem 2.4).
We emphasise that this is not a statement about floating-point arithmetic per se. The abstraction intentionally excludes scale-adaptive regimes where thresholds/biases grow with fan-in or where weights shrink with dimension (e.g. $1/n$ or $1/\sqrt{n}$ normalisations), which move the analysis toward threshold-circuit classes such as $\mathrm{TC}^{0}$.
Remark 2.2 (Scope of the $\mathrm{AC}^{0}$ abstraction).
Theorem 2.4 formalises a fixed-latency fixed-scale quantised regime in which each unit uses parameters drawn from a bounded-range grid $\mathcal{Q}_{W,D}$ with $W,D$ independent of the input length. This yields the constant-witness property exploited in Lemma B.1.
The abstraction is deliberately conservative: it does not model architectures that rely on fan-in dependent re-scaling, numerically large signed biases/thresholds that grow with dimension, or deep computations whose depth scales with $n$. Allowing such scale-adaptive behaviour leads naturally to richer threshold-circuit models (e.g. $\mathrm{TC}^{0}$), in which parity-like computations can become representable.
Remark 2.3 (Local vs. global parity).
When we say “parity is hard” we mean global parity across the full Boolean input of length $n=NB$. In contrast, parity on $O(1)$ bits is trivial in $\mathrm{AC}^{0}$ (hardwire the truth table), and more generally XOR of a constant number of $B$-bit blocks is $\mathrm{AC}^{0}$-computable. This local/global separation is exactly what underlies the two-frame bypass via the nimber difference $\Delta(P,P^{\prime})$.
Property of deployed inference Captured here Not captured here
Fixed latency / bounded depth constant depth $L$ (and window $T$) depth growing with $n$
Fixed-scale quantisation (bounded range) weights/thresholds in fixed grid $\mathcal{Q}_{W,D}$ fan-in dependent scaling or thresholds growing with $n$
Parallel bit operations unbounded fan-in AND/OR/NOT exact global counting / exact parity on unbounded inputs
Signed effects via input negations (Lemma B.1) large signed dot-product accumulation
Notation convention (input length).
We write $N$ for the number of heaps and $B$ (or $k$ in experiments) for the bit-width used to encode each heap size. The resulting Boolean input length is $n:=NB$. When we discuss $\mathrm{AC}^{0}$ circuit size/depth, $n$ always refers to this Boolean input length.
Circuit complexity primer.
$\mathrm{AC}^{0}$ is the class of Boolean functions computable by constant-depth, polynomial-size circuits with unbounded fan-in AND/OR/NOT gates. A classical result is that $\mathrm{PARITY}\notin\mathrm{AC}^{0}$. In contrast, $\mathrm{TC}^{0}$ augments $\mathrm{AC}^{0}$ with unbounded fan-in threshold gates and can compute parity. Our focus here is a fixed-scale quantised regime in which thresholds do not scale with fan-in; in that regime, each threshold unit admits a constant-witness DNF expansion and the overall computation collapses to $\mathrm{AC}^{0}$ (Theorem 2.4).
@@ 2.2 Simulation by circuits
Theorem 2.4 (Simulation of fixed-scale quantised neural models by $\mathrm{AC}^{0}$).
@@
    (Fixed-scale quantisation) all weights and thresholds lie in $\mathcal{Q}_{W,D}$.
@@
3 Related Work
Circuit complexity and neural networks.
The correspondence between neural computation and circuit classes is foundational. Parity is outside $\mathrm{AC}^{0}$ [5, 7], while constant-depth threshold circuits ($\mathrm{TC}^{0}$) are more expressive. Recent analyses connect modern attention architectures to this lens [10, 6]. In particular, Merrill and Sabharwal show that parity-like behaviour remains difficult for standard transformer computation without additional structured reasoning channels [10]. Our contribution differs in focus: we model the control policy in RL in a fixed-latency, fixed-scale quantised inference regime and study the consequences for planning.
Quantized inference and deployment constraints.
Our fixed-scale assumption is aligned with practical low-bit deployment regimes. Integer-arithmetic-only inference with fixed scaling and bounded dynamic range is standard in efficient neural deployment pipelines [13, 14]. We use this as a conservative theoretical abstraction: not to model every floating-point detail, but to capture a regime where threshold scales do not grow with input dimension.
Reinforcement learning in combinatorial games.
While RL has mastered domains such as Go and chess, impartial games remain a useful stress test because optimality is governed by exact algebraic invariants rather than approximate pattern matching. Friedman [4] explicitly proposed NIM as an AI challenge for this reason. Zhou and Riis [31, 32] reported empirical failures of standard DQN/PPO-style pipelines on NIM. Our results provide a formal impossibility account that explains those observations in a fixed-scale quantised, fixed-latency regime.
History and memory in RL.
Using short history windows is standard in partially observable settings, where stacked frames recover hidden variables such as velocity [12]. Our point is different and perhaps surprising: even in Markovian, fully observable games—where the current state is informationally sufficient and the past is theoretically irrelevant—history can reduce computational complexity by exposing local differences that would otherwise require expensive search or unbounded-depth inference to extract from the current state alone. In our setting, two consecutive frames turn a parity-heavy global decision into an $\mathrm{AC}^{0}$-computable local restoration rule: the previous frame effectively donates, for free, the computational work that search would need to perform.
@@ 4 Impartial games and NIM
We briefly recall standard definitions; see [2] for extensive background.
@@ 4.2 NIM
A NIM position consists of $h$ heaps with sizes $x_{1},\dots,x_{h}\in\mathbb{N}_{0}$. A legal move reduces exactly one heap: choose $i$ with $x_{i}>0$ and replace $x_{i}$ by some $x_{i}^{\prime}<x_{i}$.
@@ 4.3 Sprague–Grundy theory and nim-sum
Why NIM is the canonical barrier.
The Sprague–Grundy theorem implies that mastering any impartial game reduces to two tasks: (1) computing the Grundy values (nimbers) of sub-components, and (2) computing their nim-sum. Our impossibility results target the second step. By proving that fixed-scale networks cannot compute the nim-sum (Theorem 5.2), we establish a fundamental barrier for the entire class of impartial games: no matter how easily an agent might perceive the local structure of a game (e.g., in Kayles or Dawson’s Kayles), it cannot aggregate these local values into a global decision whenever the aggregation logic relies on parity. Thus, NIM is not merely a specific game, but the isolate computation required to solve combinatorial sums of games.
@@ 5.1 Mastery notions
We distinguish three notions of mastery:
Definition 5.1 (Strong, weak, and restoration mastery).
An agent achieves strong mastery of NIM if, for every reachable position $p$ with the agent to move, it plays optimally (i.e., implements a winning strategy whenever one exists).
It achieves weak mastery relative to an initial position $p_{0}$ if it plays optimally on every position that is reachable from $p_{0}$ under some sequence of legal moves, against any opponent strategy.
It achieves restoration mastery if, whenever the opponent moves from a $P$-position $Q$ with $\operatorname{nim}(Q)=0$ to a position $Q^{\prime}$, the agent responds with a move $Q^{\prime}\to R$ such that $\operatorname{nim}(R)=0$.
We use standard impartial-game terminology: a $P$-position is losing for the player to move, and an $N$-position is winning. For NIM, Theorem 4.4 says that $P$-positions are exactly the positions with nim-sum 0.
@@ 5.2 Single-frame impossibility results
Theorem 5.2 (Single-frame strong mastery requires parity).
Consider the restricted family of NIM positions $P(\mathbf{b})=(2,b_{1},\dots,b_{m})$ where $b_{i}\in\{0,1\}$. Any single-frame agent that achieves strong mastery must compute the parity $b_{1}\oplus\cdots\oplus b_{m}$. Since $\mathrm{PARITY}\notin\mathrm{AC}^{0}$, no polynomial-size, constant-depth fixed-scale quantised network can strongly master NIM.
Proof.
Let $s=\operatorname{nim}(2,b_{1},\dots,b_{m})=2\oplus(b_{1}\oplus\cdots\oplus b_{m})$. The most significant bit of $s$ is $2$ (the $2^{1}$ bit), and among the heaps only the first heap has that bit set. Therefore, any move that makes the nim-sum zero must change the first heap. The unique choice that makes the new nim-sum zero is
$x_{1}^{\prime}\;=\;2\oplus s\;=\;b_{1}\oplus\cdots\oplus b_{m}\in\{0,1\},$
so the winning move is uniquely determined by the parity of $(b_{1},\dots,b_{m})$.
Moreover, the Boolean input length $n$ of this family satisfies $n=\Theta(m)$ under any fixed-width binary encoding of the heaps, so this yields a standard $\mathrm{AC}^{0}$ hardness reduction from PARITY. ∎